Hello all having trouble figuring this problem out. Using only 1 variable and writing an inequality to come up with a solution how would I do this problem.
The length of a rectangular boat storage shed must be 4 meters more than the width . The perimeter must be at least 120 meters. What widths of the shed will make the perimeter at least 120 meters?

Answer by nerdybill(7384)

The length of a rectangular boat storage shed must be 4 meters more than the width . The perimeter must be at least 120 meters. What widths of the shed will make the perimeter at least 120 meters?
.
Let w = width
then
w+4 = length
.
2(w + w+4) >= 120
2(2w+4) >= 120
2w+4 >= 60
w+2 >= 30
w >= 28 meters
width must be "greater than or equal" to 28 meters